OpenStudy (anonymous):

Let f(x) = (1-4x) / (4-x) Find f^-1 Which is actually the inverse

OpenStudy (anonymous):

Okay....so the inverse of a function is a function such that \[f(f ^{-1}(x))=1\] Right?

OpenStudy (anonymous):

I thought it would be such that f(f^-1(x)) = x Im probably wrong...

OpenStudy (anonymous):

No you are right. :) It's been a little while since I've done inverses.

OpenStudy (anonymous):

Okay, to find the inverse we can solve y=f(x) for x: y=(1-4x)/(4-x) (4-x)*y=1-4x 4y-xy+4x-1=0 4y-1-xy+4x=0 4y-1-x(y+4)=0 -x(y+4)=1-4y -x=(1-4y)/(y+4) x=(4y-1)/y+4 = f^(-1)(x) Check my arithmetic and algebra cause I get really sloppy!!

OpenStudy (anonymous):

Excellent!